# Ming-Flash-Omni: A Sparse, Unified Architecture for Multimodal Perception and Generation

### v1

@@ 3.3 Video Data
Video streaming conversation constitutes a fundamental capability for MLLMs in video understanding. However, acquiring large-scale streaming multi-turn conversation data is prohibitively expensive. In this work, we propose a pipeline to systematically synthesize diverse, balanced, and high-quality multi-turn conversation video datasets. We collect 8.2M videos from the internet, ranging from 90 seconds to 10 minutes, and first filter out low-quality videos with high speech density, high shot density, irregular aspect ratios, or low resolution. We then filter out low-information, incoherent, or overly simple videos using SOTA MLLMs. To ensure a balanced dataset, we use advanced embedding models (e.g., Qwen3-Embedding (Zhang et al., 2025) and M3-Embedding (Chen et al., 2024a)) to extract embeddings and then cluster the videos to suppress high-frequency data while preserving long-tail content. Finally, we use SOTA video understanding models to generate high-quality video conversations. This produces 1.2M conversation turns across 5-minute average videos, balanced across various task categories.
@@ 3.4 Text Data
For text data, we utilize corpus from Ling (LingTeam et al., 2025), M2-Omni (Guo et al., 2025), and Ming-Omni (AI et al., 2025a) to preserve and further enhance the model’s language proficiency.
@@
Table 4: Performance of Ming-Flash-Omni on Image-to-Mask Segmentation Benchmarks compared to leading models. Model types are denoted as: Vision. for vision-only models, SAM. for models equipped with an additional SAM-like segmentation head, and Unified. for unified MLLMs capable of both understanding and generation. Results with “*“ are obtained by evaluating on 500 images sampled from each dataset via the official API.
@@ 4.2 In-house Benchmarks
Video Streaming Multi-turn Benchmark. The evaluation of video streaming multi-turn dialogue capabilities requires quantifying not only the model’s understand capability but also assessing its interactive experience, including proactivity and naturalness. Previous streaming dialogue datasets, such as StreamBench (Lin et al., 2024) and OvO-Bench (Niu et al., 2025), have primarily focused on the understanding aspect while lacking a thorough evaluation of the interactive experience. To address this gap, we introduce StreamingMultiturnBench. To construct StreamingMultiturnBench, we manually selected 380 videos, carefully ensuring coverage of multiple key domains including life recording, education, TV shows, video games, and documentaries. Then we use SOTA closed-source modelfor machine annotation. Subsequently, a team of 10 human annotators revise and double-check the dialogue content to ensure it aligns with human conversational preferences. This process yielded 2,200 video question-answer pairs. During evaluation, we use advanced closed-source model, e.g. GPT-4o (OpenAI, 2025), to compare the model’s output against the human-annotated answers, scoring it on a scale of 1 to 5 across the six dimensions: accuracy, completeness, relevance, naturalness, conciseness, and proactivity. The final score is the average for each dimension. To align our metrics with other video benchmarks, we linearly scale the results to a 100-point scale. We commit to open-sourcing and publicly maintaining this benchmark to ensure reproducibility.
@@ 4.3 Quantitative Results
Audio $\rightarrow$ Text (Understanding). Our model sets a new state-of-the-art (SOTA) on all 12 sub-tasks of the ContextASR-Bench (Table 6), underscoring its superior ability to leverage context—a vital skill for real-world applications like multi-turn dialogue and hotword enhancement. It also exhibits highly competitive performance across various ASR benchmarks, with notable strengths in dialect recognition (Table 8). In audio question answering, Ming-Flash-Omni surpasses all open-source audio-centric and other Omni models, with the exception of Qwen3-Omni-Flash-Instruct(Xu et al., 2025b) (Table 8). Taken together, these findings demonstrate the robust and versatile audio understanding capabilities of Ming-Flash-Omni.
Text $\rightarrow$ Audio (Generation). As shown in Table 6, leveraging advancements in speech representation and model architecture, Ming-Flash-Omni achieves SOTA performance among open-source models on the test-zh subset of the SEED-TTS-Eval benchmark(Anastassiou et al., 2024b). Furthermore, its WER on the test-en subset is surpassed only by that of Qwen3-Omni.
Video + Audio $\rightarrow$ Text (Video Streaming Conversation). As shown in Table 9, benefiting from the introduction of high-quality and diverse streaming video multiturn data, Ming-Flash-Omni has achieved significant improvements in all dimensions compared to Ming-Lite-Omni. Ming-Flash-Omni also outperforms Qwen2.5-Omni (Xu et al., 2025a) in the dimensions of accuracy, completeness, relevance, and conciseness, providing better experience in streaming video conversation scenarios.

### v2

@@ 3.3 Video Data
Video streaming conversation constitutes a fundamental capability for MLLMs in video understanding. However, acquiring large-scale streaming multi-turn conversation data is prohibitively expensive. In this work, we propose a pipeline to systematically synthesize diverse, balanced, and high-quality multi-turn conversation video datasets. We collect 8.2M videos from the internet, ranging from 90 seconds to 10 minutes, and first filter out low-quality videos with high speech density, high shot density, irregular aspect ratios, or low resolution. We then filter out low-information, incoherent, or overly simple videos using SOTA MLLMs. To ensure a balanced dataset, we use advanced embedding models (e.g., Qwen3-Embedding (zhang2025qwen3) and M3-Embedding (chen2024bge)) to extract embeddings and then cluster the videos to suppress high-frequency data while preserving long-tail content. Finally, we use SOTA video understanding models to generate high-quality video conversations. This produces 1.2M conversation turns across 5-minute average videos, balanced across various task categories.
@@ 3.4 Text Data
For text data, we utilize corpus from Ling (ling), M2-Omni (guo2025m2), and Ming-Omni (ai2025ming) to preserve and further enhance the model’s language proficiency.
@@
Table 4: Performance of Ming-Flash-Omni on Image-to-Mask Segmentation Benchmarks compared to leading models. Model types are denoted as: Vision. for vision-only models, SAM. for models equipped with an additional SAM-like segmentation head, and Unified. for unified MLLMs capable of both understanding and generation. Results with “*” are obtained by evaluating on 500 images sampled from each dataset via the official API.
@@ 4.2 In-house Benchmarks
Video Streaming Multi-turn Benchmark. The evaluation of video streaming multi-turn dialogue capabilities requires quantifying not only the model’s understand capability but also assessing its interactive experience, including proactivity and naturalness. Previous streaming dialogue datasets, such as StreamBench (lin2024streamingbench) and OvO-Bench (niu2025ovo), have primarily focused on the understanding aspect while lacking a thorough evaluation of the interactive experience. To address this gap, we introduce StreamingMultiturnBench. To construct StreamingMultiturnBench, we manually selected 380 videos, carefully ensuring coverage of multiple key domains including life recording, education, TV shows, video games, and documentaries. Then we use SOTA closed-source modelfor machine annotation. Subsequently, a team of 10 human annotators revise and double-check the dialogue content to ensure it aligns with human conversational preferences. This process yielded 2,200 video question-answer pairs. During evaluation, we use advanced closed-source model, e.g. GPT-4o (chatgpt4o), to compare the model’s output against the human-annotated answers, scoring it on a scale of 1 to 5 across the six dimensions: accuracy, completeness, relevance, naturalness, conciseness, and proactivity. The final score is the average for each dimension. To align our metrics with other video benchmarks, we linearly scale the results to a 100-point scale. We commit to open-sourcing and publicly maintaining this benchmark to ensure reproducibility.
@@ 4.3 Quantitative Results
Audio $\rightarrow$ Text (Understanding). Our model sets a new state-of-the-art (SOTA) on all 12 sub-tasks of the ContextASR-Bench (Table 6), underscoring its superior ability to leverage context—a vital skill for real-world applications like multi-turn dialogue and hotword enhancement. It also exhibits highly competitive performance across various ASR benchmarks, with notable strengths in dialect recognition (Table 8). In audio question answering, Ming-Flash-Omni surpasses all open-source audio-centric and other Omni models, with the exception of Qwen3-Omni-Flash-Instruct(xu2025qwen3omnitechnicalreport) (Table 8). Taken together, these findings demonstrate the robust and versatile audio understanding capabilities of Ming-Flash-Omni.
Text $\rightarrow$ Audio (Generation). As shown in Table 6, leveraging advancements in speech representation and model architecture, Ming-Flash-Omni achieves SOTA performance among open-source models on the test-zh subset of the SEED-TTS-Eval benchmark(anastassiou2024seed). Furthermore, its WER on the test-en subset is surpassed only by that of Qwen3-Omni.
Video + Audio $\rightarrow$ Text (Video Streaming Conversation). As shown in Table 9, benefiting from the introduction of high-quality and diverse streaming video multiturn data, Ming-Flash-Omni has achieved significant improvements in all dimensions compared to Ming-Lite-Omni. Ming-Flash-Omni also outperforms Qwen2.5-Omni (xu2025qwen2) in the dimensions of accuracy, completeness, relevance, and conciseness, providing better experience in streaming video conversation scenarios.
@@ 6 Contributors
[…]
7 Public Benchmarks
Image $\rightarrow$ Text (Understanding). Our evaluation of the image-to-text understanding capabilities primarily encompasses the following six tasks: 1) general image understanding capabilities evaluated on MMStar (chen2024we), AI2D (kembhavi2016diagram), HallusionBench (Guan_2024_hallusionbench), CV-Bench (tong2024cambrian1), MathVista (lu2024mathvista), and CRPE (wang2024crpe). 2) OCR capabilities evaluated on ChartQA (masry-etal-2022-chartqa), DocVQA (docvqa), OCRBench (Liu_2024), and TextVQA-VAL (Eval_textvqa). 3) multi-image capabilities evaluated on MMTBench (ying2024mmtbench), MuirBench (wang2024muirbench), and LLaVA-interleave Bench (li2024llava_interleave_bench). 4) Complex instruction capabilities evaluated on MIA-Bench (qian2024mia). 5) video understanding capabilities evaluated on MVBench (li2024mvbench), VideoMME (fu2024videomme), and LongVideoBench (wu2024longvideobench).
Text $\rightarrow$ Image (Generation). We incorporate text-to-image generation capabilities to enable our MLLM with unified perception-generation abilities, which are evaluated on GenEval (ghosh2024geneval), DPG-Bench (hu2024ella), and FID.
Image $\rightarrow$ Image (Editing). Our evaluation of image-to-image editing capabilities is conducted on the GEdit-Bench benchmark (liu2025step1x).
Image $\rightarrow$ Image (Segmentation). We evaluate the segmentation capability of our MLLM on the standard referring expression segmentation (RES) benchmarks RefCOCO/+ (refcoco) and RefCOCOg (refcocog).
Audio $\rightarrow$ Text (Understanding). Our evaluation of the audio-to-text understanding capabilities mainly includes the following three tasks: 1) Fundamental audio understanding capabilities evaluated on a broad range of public benchmarks, including public Chinese benchmarks like Aishell1 (AISHELL1) and Wenetspeech (WenetSpeech), and public English benchmarks like Librispeech (Librispeech) and Voxpopuli (wang2021voxpopuli). And 2) audio question-answering capabilities evaluated on various benchmarks across five specific tasks, such as AlpacaEval and CommonEval from VoiceBench (chen2024voicebenchbenchmarkingllmbasedvoice) for open-ended QA tasks, and SD-QA for knowledge-based QA tasks. Finally 3) evaluates the model’s ability to utilize context on ContextASR-Bench(wang2025contextasrbenchmassivecontextualspeech).
Text $\rightarrow$ Audio (Generation). We incorporate text-to-audio generation capabilities to enable our MLLM with unified audio perception-generation abilities, which are evaluated on Seed-TTS-Eval (Seed_TTS).
Video $\rightarrow$ Text (Understanding). Our evaluation of the video-to-text understanding capabilities contains the following four benchmarks: MVBench (li2024mvbench), VideoMME (fu2024video) and LongVideoBench (wu2024longvideobench).
8 The prompt used in data generation
This section presents the prompt used for data generation in Sec. 3.
Prompt for portrait preservation data Based on the persons age and gender in the image, generate a detailed description of a realistic life scene photo. Requirements: 1. Must include: clothing, location, action, style. 2. Strictly within 50 words. 3. Change the scene (real scenes: indoor, outdoor, residential areas, homes, companies, kitchens, parks, etc.) and character costumes to ensure that even similar images have different descriptions. 4. Clothing must match age and gender, vary descriptions. 5. Scenario details must include at least 3 realistic elements, be vivid and lifelike. Output only the description in English, no additional text. Example: A young woman with medium-length dark hair tied back in a neat ponytail stands outdoors near a park bench, dressed in a crisp white button-up shirt under a fitted black blazer. She wears a subtle headband, her brow knitted as she runs her fingers through her hair, her lips slightly downturned and eyes glistening with unshed tears. The scene captures her from a distant angle, with trees and a blurred pathway in the soft-focus background, suggesting contemplation or distress amidst nature.
Prompt for text generation data Describe according to the following requirements: 1. Randomly select a theme from {theme}, and generate text content in Chinese, English, and numbers, with 3-5 characters. 2. Use your imagination; the theme is not fixed. Keep the description under 100 characters. 3. Generate a suitable image description based on the text content. 4. Refer to the {text style} and choose appropriate font, color, and layout. Output format: Description, Text: "content", Font style, Font color, and Image layout Examples: In the image, there is a gym scene, Text: "Power 3.0", font style is broken art font with cracks and shattering effects, font color is metallic gray with dark red gradient, text layout is centered-right, background includes dumbbells, treadmills, and reflective glass walls.
Prompt for video multi-turn conversation data # Role: Expert Video Analyst You are an expert video analyst. Your task is to analyze the provided video and output a structured JSON object containing your analysis. You must adhere strictly to the format and rules described below. # Instructions: Analyze the video and generate a single JSON object with the following keys. Your response should ONLY be the JSON object, enclosed in a single markdown code block (‘‘‘json ... ‘‘‘). Do not include any other text, explanations, or introductory phrases. # JSON Fields Definition: 1. "category"‘: (String) Select ONLY ONE category from the following list that best describes the main subject of the video. If the video category is not among the listed below, output "Others". * **List of 34 Categories**: [ "Others", "LifeRecord-TravelLog", "LifeRecord-DailyLife", "LifeRecord-HouseTour", "LifeRecord-Reaction", "LifeRecord-AnimalPet", "LifeRecord-Cooking", "LifeRecord-Fashion", "LifeRecord-Workout", "Education-Lecture", "Education-Finance", "Education-Multilingual", "Education-Handifraft", "Education-Science", "Education-Art", "Education-OnlineTutorial", "TVShow-TVSeries", "TVShow-News", "TVShow-TalkShow", "TVShow-Celebration", "TVShow-CommentaryProgram", "Competition-Football", "Competition-Athletics", "Competition-Basketball", "Competition-Snooker", "Competition-Boxing", "Competition-Car", "VideoGames-Sandbox", "VideoGames-OpenWorld", "Documentary-Nature", "Documentary-Science", "Documentary-Culture", "Documentary-Kids", "Movie-Comedy", "Movie-Adventure"] 2. "caption"‘: (String) A concise but descriptive summary of the videos content in English, not exceeding 30 words. Describe what is happening, who is involved, and the main subject. 3. "content_score"‘: (Integer) An integer from 1 to 10. This score evaluates the videos potential as a prompt for a synthetic AI-user conversation. * **Score 1-3 (Low)**: The video content is sparse, repetitive, or features a single persons monologue with little interaction or environmental detail. Its difficult to ask questions or start a conversation based on it. * **Score 4-7 (Medium)**: The video has some interesting elements but may lack depth or variety. It can support a few questions but may not lead to an extended, rich conversation. * **Score 8-10 (High)**: The video is rich in content, detail, and action. It shows a process, an interaction, or a complex scene that naturally invites questions and allows for a deep, extended conversation (e.g., a detailed cooking tutorial, a complex assembly process, a travel vlog with multiple activities). 4. "fov"‘: (Integer) Field of View. * **1**: The video is shot from a first-person perspective (FPV), where the camera acts as the viewers eyes. * **0**: The video is shot from a third-person or static perspective.
5. "task_complexity"‘: (Integer) An integer from 1 to 10, representing the complexity of the primary task shown in the video. If no specific task is shown, rate the complexity of the main activity. * **Score 1-3 (Low)**: Simple, everyday actions that require minimal skill (e.g., opening a bottle, pouring water, petting a cat). * **Score 4-7 (Medium)**: Tasks that require some skill, knowledge, or multiple steps (e.g., cooking a simple dish, assembling IKEA furniture, basic makeup application). * **Score 8-10 (High)**: Highly complex, specialized, or professional tasks that require significant expertise, precision, or effort (e.g., performing surgery, building a car engine, professional programming, playing a complex musical piece). # Required Output Format: ‘‘‘json { "category": "...", "caption": "...", "content_score": ..., "fov": ..., "task_complexity": ... } ‘‘‘